Calculate the volume in liters of a 0.090mM potassium permanganate solution that contains 9.00mol of potassium permanganate KMnO4. Round your answer to 2 significant digits.
Answer #1

Answer

1.0×105 L

Explanation

Molarity is defined as the number of moles of solute per liter of solution

1mM =0.001M

0.090mM = 0.000090M

0.000090M means 0.000090moles of KMnO4 in 1liter

Therefore

Volume of 0.090mM KMnO4 solution containing 9.00moles of KMnO4 = (1000ml/0.000090mol)×9.00mol = 1.0 ×108ml = 1.0×105 L